Big TimeBig Time Investor Group is opening an office in​ Portland, Oregon. Fixed monthly costs are office rent ​($$8,900​), depreciation on office furniture left parenthesis $($1,700)​, utilities ​($2,500​), special telephone lines ​($1,600​), a connection with an online brokerage service ​($2,600​), and the salary of a financial planner ​($17,700​). Variable costs include payments to the financial planner ​(88​% of​ revenue), advertising left parenthesis (13% of​ revenue), supplies and postage ​(33​% of​ revenue), and usage fees for the telephone lines and computerized brokerage service left parenthesis 6 %(6% of​ revenue).

1.

Use the contribution margin ratio approach to compute

Big Time​'s breakeven revenue in dollars. If the average trade leads to  $1,000
in revenue for Big Time​, how many trades must be made to break​ even?

2.

Use the equation approach to compute the dollar revenues needed to earn a monthly target profit of $12,600.

3.

Graph

Big Time​'s CVP relationships. Assume that an average trade leads to $1,000

in revenue for Big Time. Show the breakeven​ point, the sales revenue​ line, the fixed cost​ line, the total cost​ line, the operating loss​ area, the operating income​ area, and the sales in units​ (trades) and dollars when monthly operating income of $12,600 is earned.

4.

Suppose that the average revenue

Big Time earns increases to $2,000 per trade. Compute the new breakeven point in trades. How does this affect the breakeven​ point?

​(Round your answers to the nearest whole​ number.)
Answer #1

Q - 1

Contribution margin % = 1 - Variable cost as % of sales = 1 - financial planner cost ​- advertising - supplies and postage ​- usage fees for the telephone lines and computerized brokerage service = 1 - 8% - 13% - 3% - 6% = 70%

Fixed costs = Office rent ​+ depreciation on office furniture left parenthesis + utilities ​+ special telephone lines ​+ a connection with an online brokerage service ​+ and the salary of a financial planner = 8,900​ + 1,700 + 2,500 + 1,600 + 2,600 + 17,700 = $ 35,000

Big Time​'s monthly breakeven revenue in dollars = Monthly fixed costs / Contribution margin % = 35,000 / 70% = $ 50,000

Trades needed to breakeven monthly = 50,000 / 1,000 = 50

Q - 2

Equation for profit:

Profit = (1 - Variable cost as % of sales) x Revenue - Fixed costs

12,600 = (1 - 30%) x Revenue - 35,000

Hence, Revenue = (12,600 + 35,000) / (1 - 30%) = $ 68,000

Q - 3

The CVP relationship can be plotted using the table below:

Number of trades Sales Fixed cost Total Cost
n =1000 x n =Sales x (1 - 30%) - Fixed cost
                -                     -           35,000                                          35,000
               10          10,000         35,000                                          38,000
               20          20,000         35,000                                          41,000
               30          30,000         35,000                                          44,000
               40          40,000         35,000                                          47,000
               50          50,000         35,000                                          50,000
               60          60,000         35,000                                          53,000
               70          70,000         35,000                                          56,000
               80          80,000         35,000                                          59,000
               90          90,000         35,000                                          62,000
            100       100,000         35,000                                          65,000

The graph is shown below:

CVP Plot 120000 100000 80000 60000 A 40000 20000 0 60 0 10 20 30 40 50 70 80 an 100 110 Number of trades Sales Fixed cost Tot Point A is the break even point. Area shaded in red line is the area of operating losses. Area shaded in green line is area of operating profit.

Q - 4

Average revenue per trade = $ 2,000 which is twice the average revenue per trade initially. Hence the break even number of trades will reduce to half. New break even trade number = 50,000 / 2,000 = 25